# Erase then Rectify: A Training-Free Parameter Editing Approach for Cost-Effective Graph Unlearning

### v1

@@ Introduction
Machine unlearning has garnered considerable attention in recent years due to its ability to remove the influence of specific subsets of data from a well-trained machine learning model (Nguyen et al. 2022). This process is particularly important for forgetting sensitive, mislabeled, or obsoleted information. Typically, the subset of data to be unlearned is much smaller than the entire training dataset. As a result, one primary goal of machine unlearning is to efficiently eliminate the impact of these unlearned samples while preserving the model’s predictive power, thereby avoiding the costly process of retraining the model from scratch on the remaining data (Xu et al. 2024).
@@
In the graph domain, there is also a strong demand to forget specific information. For example, users in a social network may request the removal of their personal data. In particular, Graph Neural Networks (GNNs) has become a common routine for processing and analyzing graph-structured data, leveraging message propagation and neighborhood aggregation mechanisms (Zhou et al. 2020). However, the interconnected nature of graphs determines that unlearned samples can significantly impact their neighbors through message propagation, making conventional machine unlearning methods unapplicable (Cheng et al. 2023). Consequently, graph unlearning has emerged as a specialized machine unlearning task with the goal of removing unlearned samples and mitigating their influence on other intercorrelated nodes (Said et al. 2023).
@@
Recently, several approaches have been proposed to address the challenges of graph unlearning. For instance, GIF (Wu et al. 2023a) and CGU (Chien, Pan, and Milenkovic 2023) directly approximate the parameter changes induced by data removal and update the affected parameters. Such methods primarily focus on forgetting unlearned samples while overlooking the predictive capability on remaining data, which can significantly compromise model utility (Li et al. 2024). To resolve this problem, some recent studies (Li et al. 2024; Wang, Huai, and Wang 2023) have simultaneously considered both predictive and unlearning objectives when an unlearning request is received. However, they typically necessitate additional training on the remaining data, which leads to significant computational overhead, particularly with large graphs, even if only a small number of unlearned samples are involved. The substantial computational resources required constrain the efficiency and scalability of these methods, thereby limiting their practicality in real-world applications.
@@
In this work, we focus on developing an efficient and scalable graph unlearning method that preserves the utility of GNNs. However, achieving this goal is non-trivial. First, the interconnected nature of graph data means that unlearned samples can significantly affect their neighbors through message propagation, making it challenging to efficiently remove both the unlearned samples and their influence on other nodes. Second, the remaining dataset after unlearning is typically much larger than the unlearned subset, which can lead to substantial computational overhead when accessing the remaining data. Balancing the need to maintain model utility with the demand for low computational overhead presents another significant challenge.
@@
To tackle the above challenges, we propose Erase then Rectify (ETR), a training-free, two-stage approach for cost-effective graph unlearning. Specifically, we first establish a theoretical foundation demonstrating that masking parameters critical to unlearned samples enables effective graph unlearning. Building on this theory, in the Erase stage, we propose a neighborhood-aware parameter editing strategy to effectively eliminate the impact of unlearned samples and their cascading effects on the entire graph, while minimizing the impact on model utility. In the Rectify stage, we introduce a subgraph-based gradient approximation method to efficiently estimate the gradient of the unlearned model on the remaining data, which is then used to further enhance the model performance. Notably, ETR achieves graph unlearning without requiring explicit model training or access to the entire training dataset, which not only avoids substantial computational overhead but also ensures data privacy and security. Overall, the main contributions are as follows:
@@
Graph unlearning methods can be categorized into exact and approximate approaches, aiming to obtain an unlearned model that is exactly or approximately equivalent to retraining from scratch (Said et al. 2023). Current efforts primarily focus on approximate graph unlearning. For example, GraphEraser (Chen et al. 2022) and GUIDE (Wang, Huai, and Wang 2023) follow the SISA (Bourtoule et al. 2021) paradigm, proposing graph partitioning algorithms to divide the training set into multiple shards, with a separate model trained for each shard. Besides, some approaches (Wu et al. 2023a, b; Chen et al. 2023; Chien, Pan, and Milenkovic 2023) leverage influence functions to approximate the impact of unlearning requests on model parameters. GNNDelete (Cheng et al. 2023) freezes model parameters while training additional parameters to achieve graph unlearning, and MEGU (Li et al. 2024) simultaneously trains the model’s prediction and unlearning objectives. Despite the effectiveness, these methods still face limitations when scaling to large graphs.
@@
Several training-free methods have been proposed for efficient machine unlearning. For example, Fisher Forgetting (Golatkar, Achille, and Soatto 2020a) and NTK (Golatkar, Achille, and Soatto 2020b) are weight scrubbing methods that add noise to parameters informative for unlearned samples. However, Fisher Forgetting is computationally expensive and sacrifices the model’s predictive performance (Tarun et al. 2023), while NTK relies on additional models to achieve unlearning. The most relevant work to ours is SSD (Foster, Schoepf, and Brintrup 2024), which achieves unlearning by dampening parameters crucial for the unlearned dataset but not for the remaining dataset. Unlike SSD, we account for the impact of unlearned samples on other samples through message propagation. Additionally, we propose adaptively selecting hyperparameters, enhancing the method’s versatility across different datasets. Furthermore, we introduce the Rectify strategy to improve the model performance on the remaining dataset.
@@
In this paper, we employ $\mathcal{G}=\left(\mathcal{V},\mathcal{E},\mathcal{X}\right)$ to denote a graph comprising $\left|\mathcal{V}\right|$ nodes and $\left|\mathcal{E}\right|$ edges. Each node $\mathbf{v}_{i}\in\mathcal{V}$ has a $d$-dimensional feature vector $\mathbf{x}_{i}\in\mathcal{X}$. A major category of GNNs is message-passing neural networks (MPNNs) (Gilmer et al. 2017), including models such as GCN (Kipf and Welling 2017) and GAT (Velickovic et al. 2018). MPNNs follow the “propagate-transform” paradigm: in each layer of the network, they propagate and aggregate features from neighboring nodes, and then transform these aggregated features to update each node’s representation (Gilmer et al. 2017). Following previous works (Chen et al. 2022; Said et al. 2023), we focus on the task of node classification in this study. We denote the training dataset as $D$, the unlearned dataset as $D_{f}$, the remaining dataset as $D_{r}$ (i.e., $D\setminus D_{f}$), and the $k$-hop neighborhood of $D_{f}$ as $D_{k}$.
@@
The FIM utilizes the second-order derivative of the loss function to quantify the sensitivity of each parameter, thereby providing a measure of its importance with respect to the input samples (Guo et al. 2020). Specifically, for a distribution $p(y|x,w)$, the FIM and its first-order derivative property (Kay 1993) are given as follows:
@@
Notably, obtaining the FIM is computationally expensive. A common approach is to use its diagonal $diag(F)$ (Kirkpatrick et al. 2017), which can be computed efficiently using the first-order derivative. The $i$-th element of the diagonal of the FIM calculated over the dataset $D$ is denoted as $F_{D,ii}$, representing the importance of the $i$-th parameter with respect to the dataset $D$.
@@
Graph unlearning can be categorized into the following three tasks: Node Unlearning: $\Delta\mathcal{G}=\left(\Delta\mathcal{V},\Delta\mathcal{E},\Delta\mathcal{X}\right)$, where $\Delta\mathcal{V}$ represents the unlearned nodes, $\Delta\mathcal{E}$ denotes the edges connecting to $\Delta\mathcal{V}$, and $\Delta\mathcal{X}$ represents the features of nodes in $\Delta\mathcal{V}$. Edge Unlearning: $\Delta\mathcal{G}=\left(\emptyset,\Delta\mathcal{E},\emptyset\right)$, where $\Delta\mathcal{E}$ represents the set of edges to be unlearned. Feature Unlearning: $\Delta\mathcal{G}=\left(\emptyset,\emptyset,\Delta\mathcal{X}\right)$, where $\Delta\mathcal{X}$ represents the unlearned features, and will be padded with zeros. We will use the node unlearning task to illustrate the proposed method, and describe how the proposed method applies to edge unlearning and feature unlearning tasks in the Appendix.
@@
To address the above limitations, we propose the Erase strategy as follows:
@@
Note the choice of hyperparameters may vary depending on the dataset. To facilitate hyperparameter selection, we adaptively select them for different datasets. Specifically, $\gamma$ is set as the top $m$value of $\frac{F_{D_{f}}}{F_{D}}$, and $\eta$ is set as the top $m$value of $\frac{F_{D_{f}}F_{D_{k}}}{{F_{D}}^{2}}$. This ensures that $m$of the parameters satisfy $F_{D_{f},jj}>\gamma F_{D,jj}$ and $m$satisfy $F_{D_{f},jj}F_{D_{k},jj}>{\eta F_{D,jj}}^{2}$. Furthermore, we set $a=\gamma$ and $b=\eta$ to balance the magnitude between $\frac{F_{D,jj}}{F_{D_{f},jj}}$ and $\frac{{F_{D_{j}j}}^{2}}{F_{D_{f},jj}F_{D_{k},jj}}$.
@@
Backbone Method PubMed CiteSeer Cora CS Physics
@@
For edge and feature unlearning, we can forget the impact of the unlearned edges or features and enhance model performance in a similar manner. The pseudo-code for edge and feature unlearning is provided in the Appendix.
@@
We compare ETR with various baselines, including Retrain, GraphEraser-BEKM, GraphEraser-BLPA (Chen et al. 2022), GIF (Wu et al. 2023a), GUIDE-SR, GUIDE-Fast (Wang, Huai, and Wang 2023), and MEGU (Li et al. 2024). Detailed descriptions of these baselines are provided in the Appendix.
@@
Following GIF, we partition each graph into a training subgraph comprising training nodes and a test subgraph containing the remaining nodes. For PubMed, CiteSeer, Cora, CS, and Physics, we randomly allocate 90% of the nodes to the training set. For ogbn-arxiv and ogbn-products, we use the splits provided in (Hu et al. 2020). The unlearning ratio is set to 5% for all tasks. We use two-layer GCN and GAT models with a hidden state dimension of 256, training for 100 epochs. All experiments are conducted on a single Nvidia A40 GPU. Each experiment is run ten times, and we report the average value and standard deviation.
@@
Backbone Method PubMed CiteSeer Cora CS Physics
GCN Retrain 1.57 1.52 1.3 6.13 14.62
BEKM 0.8 0.6 0.45 5.87 14.87
@@
GAT Retrain 1.77 1.52 1.3 6.17 15.28
BEKM 0.78 0.6 0.45 5.87 14.86
BLPA 0.48 0.48 0.4 1.6 3.53
GIF 2.88 0.8 0.72 5.89 14.3
SR 1.52 1.23 1.2 2.2 5.21
Fast 1.52 1.23 1.2 2.41 5.32
@@
We investigate how different values of $m$ affect the performance of the Erase strategy, with results shown in Figure 3. It can be observed that increasing $m$ generally improves model performance, but performance declines if $m$ becomes too large. This is because larger $m$ values help the model forget unlearned samples effectively, while excessively large values can also lead to the loss of essential knowledge. Typically, $m$ values between 8‰ and 16‰ provide the best results for the Erase strategy. Notably, the model performs better compared to $m=0$, confirming that the Erase strategy effectively forgets unlearned samples and supports Theorem 1.
@@
In this section, we investigate the effectiveness of different strategies in ETR. The Erase is crucial for model forgetting, while the Rectify is critical for maintaining model utility. The results in Figure 3 indicate that increasing m improves performance compared to $m=0$, demonstrating the effectiveness of the Erase strategy in forgetting the unlearned samples. The results in Figure 3 show that increasing $\lambda$ also improves performance compared to $\lambda=0$, indicating that the Rectify strategy effectively enhances the model performance on the remaining dataset.
@@ Conclusion
In this paper, we investigate cost-effective graph unlearning and propose a training-free approach ETR to remove the influence of unlearned samples while maintaining the model performance on the remaining dataset. Specifically, in the Erase stage, we utilize the Fisher Information Matrix to identify parameters crucial for unlearned samples and their influence on other inter-connected samples. These parameters are then modified to forget the unlearned samples and their effects on other samples through message propagation. In the Rectify stage, we use the induced graph of the unlearned samples to efficiently approximate the model’s gradient on the remaining dataset. This gradient is then used to enhance the model performance on the remaining data. Extensive experiments on seven public datasets demonstrate that the proposed ETR method achieves an excellent trade-off between efficiency and effectiveness.
@@ Appendix C Baselines
Appendix D ETR for Edge Unlearning and Feature Unlearning
$\left|D_{i}\right|\nabla_{\widehat{\omega}_{r}}L_{D_{i}})$ (36)
Appendix E Additional Experimental Results

### v2

@@ Introduction
Machine unlearning has garnered attention for its ability to remove the influence of specific data subsets from a well-trained machine learning model (Nguyen et al. 2022), which is crucial for forgetting sensitive, mislabeled, or outdated information (Liu and Tsang 2017). Typically, the subset of data to be unlearned is much smaller than the entire training dataset. As a result, one primary goal of machine unlearning is to efficiently eliminate the impact of these unlearned samples while preserving the model’s predictive power, thereby avoiding the costly process of retraining the model from scratch on the remaining data (Xu et al. 2024).
@@
In the graph domain, there is also a strong demand to forget specific information, such as users requesting the removal of personal data in social networks. In particular, Graph Neural Networks (GNNs) are widely used to process graph-structured data through message propagation and neighborhood aggregation (Zhou et al. 2020). However, due to the interconnected nature of graphs, unlearned samples can significantly affect their neighbors, making conventional machine unlearning methods unapplicable (Cheng et al. 2023). As a result, graph unlearning has emerged as a specialized task aimed at removing unlearned samples and mitigating their impact on other nodes (Said et al. 2023).
@@
Recently, several graph unlearning approaches have been proposed. For instance, GIF (Wu et al. 2023a) and CGU (Chien, Pan, and Milenkovic 2023) approximate parameter changes caused by data removal and update the affected parameters. However, these methods primarily focus on forgetting unlearned samples while overlooking the predictive capability on the remaining data, which can significantly compromise model utility (Li et al. 2024). To address this, some recent studies (Li et al. 2024; Wang, Huai, and Wang 2023) consider both predictive and unlearning objectives when an unlearning request is received. However, they typically require additional training on the remaining data, leading to high computational overhead, particularly with large graphs, even when only a few samples are unlearned. This substantial resource demand limits the efficiency and scalability of these methods in real-world applications.
@@
In this work, we focus on developing an efficient and scalable graph unlearning method that preserves model utility. However, achieving this goal is non-trivial. First, the interconnected nature of graph data means that unlearned samples can significantly affect their neighbors through message propagation, complicating the removal of both the unlearned samples and their influence. Second, the remaining dataset is typically much larger than the unlearned subset, leading to substantial computational overhead when accessing the remaining data. Balancing model utility with low computational cost presents another significant challenge.
@@
To address these challenges, we propose Erase then Rectify (ETR), a training-free, two-stage approach for cost-effective graph unlearning. First, we establish a theoretical foundation showing that masking parameters critical to unlearned samples enables effective unlearning. Based on this, in the Erase stage, we propose a neighborhood-aware parameter editing strategy to remove the impact of unlearned samples and their cascading effects, while minimizing the impact on model utility. In the Rectify stage, we introduce a subgraph-based gradient approximation method to estimate the unlearned model’s gradient on the remaining data, further enhancing model performance. Notably, ETR achieves graph unlearning without requiring explicit model training or access to the entire training dataset, avoiding computational overhead and ensuring data privacy. Overall, the main contributions are as follows:
@@
Graph unlearning methods are divided into exact and approximate approaches, aiming to obtain a model that is exactly or approximately equivalent to retraining from scratch (Said et al. 2023). Current efforts primarily focus on approximate methods. For example, GraphEraser (Chen et al. 2022) and GUIDE (Wang, Huai, and Wang 2023) follow the SISA (Bourtoule et al. 2021) paradigm, partitioning the training set into multiple shards, with a separate model trained for each shard. Other approaches (Wu et al. 2023a, b; Chen et al. 2023; Chien, Pan, and Milenkovic 2023) use influence functions to approximate the impact of unlearning on model parameters. GNNDelete (Cheng et al. 2023) freezes model parameters while training additional parameters for unlearning. MEGU (Li et al. 2024) simultaneously trains the model’s prediction and unlearning objectives. Despite their effectiveness, these methods still face scalability issues with large graphs.
@@
Several training-free methods have been proposed for efficient machine unlearning. Fisher Forgetting (Golatkar, Achille, and Soatto 2020a) and NTK (Golatkar, Achille, and Soatto 2020b) are weight scrubbing methods that add noise to parameters informative for unlearned samples. However, Fisher Forgetting is computationally expensive and harms predictive performance (Tarun et al. 2023), while NTK requires additional models for unlearning. The closest work to ours is SSD (Foster, Schoepf, and Brintrup 2024), which dampens parameters crucial for the unlearned dataset while preserving others. Unlike SSD, we account for the impact of unlearned samples on other samples through message propagation. We also propose adaptive hyperparameter selection to improve versatility across datasets and introduce the Rectify strategy to enhance performance on remaining data.
@@
In this paper, we employ $\mathcal{G}=\left(\mathcal{V},\mathcal{E},\mathcal{X}\right)$ to denote a graph comprising $\left|\mathcal{V}\right|$ nodes and $\left|\mathcal{E}\right|$ edges. Each node $\mathbf{v}_{i}\in\mathcal{V}$ has a $d$-dimensional feature vector $\mathbf{x}_{i}\in\mathcal{X}$. A major category of GNNs is message-passing neural networks (MPNNs) (Gilmer et al. 2017), such as GCN (Kipf and Welling 2017) and GAT (Velickovic et al. 2018). MPNNs propagate and aggregate features from neighboring nodes, then transform these aggregated features to update node representations (Gilmer et al. 2017). Following previous works (Chen et al. 2022; Said et al. 2023), we focus on node classification in this study. We denote the training dataset as $D$, the unlearned dataset as $D_{f}$, the remaining dataset as $D_{r}$ (i.e., $D\setminus D_{f}$), and the $k$-hop neighborhood of $D_{f}$ as $D_{k}$.
@@
The FIM uses the second-order derivative of the loss function to quantify parameter sensitivity, providing a measure of importance with respect to input samples (Guo et al. 2020). For a distribution $p(y|x,w)$, the FIM and its first-order derivative property (Kay 1993) are given as follows:
@@
Notably, obtaining the FIM is computationally expensive. A common approach is to use its diagonal $diag(F)$ (Kirkpatrick et al. 2017), which can be computed efficiently using the first-order derivative. The $i$-th diagonal element of the FIM over dataset $D$ is denoted as $F_{D,ii}$, representing the importance of the $i$-th parameter with respect to $D$.
@@
Graph unlearning can be categorized into three tasks: Node Unlearning: $\Delta\mathcal{G}=\left(\Delta\mathcal{V},\Delta\mathcal{E},\Delta\mathcal{X}\right)$, where $\Delta\mathcal{V}$ represents unlearned nodes, $\Delta\mathcal{E}$ denotes edges connecting to $\Delta\mathcal{V}$, and $\Delta\mathcal{X}$ represents the features of nodes in $\Delta\mathcal{V}$. Edge Unlearning: $\Delta\mathcal{G}=\left(\emptyset,\Delta\mathcal{E},\emptyset\right)$, where $\Delta\mathcal{E}$ represents the set of edges to be unlearned. Feature Unlearning: $\Delta\mathcal{G}=\left(\emptyset,\emptyset,\Delta\mathcal{X}\right)$, where $\Delta\mathcal{X}$ represents the unlearned features, and will be padded with zeros. We will use the node unlearning task to illustrate the proposed method, and describe how the proposed method applies to edge unlearning and feature unlearning tasks in the Appendix.
@@
To address these, we propose the Erase strategy as:
@@
Note the choice of hyperparameters may vary depending on the dataset. To facilitate hyperparameter selection, we adaptively select them for different datasets. Specifically, $\gamma$ is set as the top $m$value of $\frac{F_{D_{f}}}{F_{D}}$, and $\eta$ is set as the top $m$value of $\frac{F_{D_{f}}F_{D_{k}}}{{F_{D}}^{2}}$. This ensures that $m$of the parameters satisfy $F_{D_{f},jj}>\gamma F_{D,jj}$ and $m$satisfy $F_{D_{f},jj}F_{D_{k},jj}>{\eta F_{D,jj}}^{2}$. Furthermore, we set $a=\gamma$ and $b=\eta$ to balance the magnitudes of $\frac{F_{D,jj}}{F_{D_{f},jj}}$ and $\frac{{F_{D_{j}j}}^{2}}{F_{D_{f},jj}F_{D_{k},jj}}$.
@@
Bone Method PubMed CiteSeer Cora CS Physics
@@
For edge and feature unlearning, we can proceed in a similar manner. The pseudo-code for edge and feature unlearning is provided in the Appendix.
@@
We compare ETR with various baselines, including Retrain, GraphEraser-BEKM, GraphEraser-BLPA (Chen et al. 2022), GIF (Wu et al. 2023a), GUIDE-SR, GUIDE-Fast (Wang, Huai, and Wang 2023), GNNDelete (Cheng et al. 2023), and MEGU (Li et al. 2024). Detailed descriptions are provided in the Appendix.
@@
Following GIF, we partition each graph into a training subgraph comprising training nodes and a test subgraph containing the remaining nodes. For PubMed, CiteSeer, Cora, CS, and Physics, we randomly allocate 90% of the nodes to the training set. For ogbn-arxiv and ogbn-products, we use the splits provided in (Hu et al. 2020). The unlearning ratio is set to 5% for all tasks. We use two-layer GCN and GAT models with a hidden state dimension of 256, training for 100 epochs. We tune the learning rates within the range of [1e-4, 1e-1] and the weight decay within [1e-7, 1e-2] for different datasets. All experiments are conducted on a single Nvidia A40 GPU. Each experiment is run ten times, and we report the average value and standard deviation.
@@
Bone Method PubMed CiteSeer Cora CS Physics
GCN Retrain 1.57 1.52 1.30 6.13 14.62
BEKM 0.80 0.60 0.45 5.87 14.87
@@
GAT Retrain 1.77 1.52 1.30 6.17 15.28
BEKM 0.78 0.60 0.45 5.87 14.86
BLPA 0.48 0.48 0.40 1.60 3.53
GIF 2.88 0.80 0.72 5.89 14.3
SR 1.52 1.23 1.20 2.20 5.21
Fast 1.52 1.23 1.20 2.41 5.32
GNNDelete 1.90 1.31 1.24 3.05 6.10
@@
We investigate the effect of different $m$ values on the Erase strategy, with results shown in Figure 3. It can be observed that increasing $m$ generally improves model performance, but performance declines if $m$ becomes too large. This is because larger $m$ values help the model forget unlearned samples effectively, while excessively large values can also lead to the loss of essential knowledge. Typically, $m$ values between 8‰ and 16‰ provide the best results for the Erase strategy. Notably, the model performs better compared to $m=0$, confirming that the Erase strategy effectively forgets unlearned samples and supports Theorem 1.
@@
This section investigates the effectiveness of different strategies in ETR. The Erase strategy is crucial for model forgetting, while Rectify is critical for maintaining model utility. Results in Figure 3 indicate that increasing $m$ improves performance compared to $m=0$, demonstrating the effectiveness of the Erase strategy in forgetting unlearned samples. Similarly, increasing $\lambda$ improves performance compared to $\lambda=0$, indicating that the Rectify strategy effectively enhances the model performance on the remaining dataset.
@@ Conclusion
In this paper, we investigate cost-effective graph unlearning and proposes the ETR method, which removes the influence of unlearned samples while preserving model performance on the remaining data. In the Erase stage, we use the Fisher Information Matrix to identify parameters crucial for unlearned samples and their impact on connected samples, then modify them to forget both the unlearned samples and their influence. In the Rectify stage, we leverage the induced graph of unlearned samples to approximate the model’s gradient on the remaining data, which is then used to enhance model performance. Extensive experiments demonstrate that ETR achieves strong performance in both efficiency and effectiveness. Future work will explore unlearning partial attributes in multi-attribute samples and partial objectives in multi-task settings (Liu and Tsang 2015; Liu, Tsang, and Müller 2017; Liu et al. 2019), where strong dependencies between different attributes and different tasks pose challenges, especially in graph-based scenarios, where there are also strong dependencies between samples.
Acknowledgements
This work was supported by the National Natural Science Foundation of China (Grant No. 62276277, 92370204), the Guangdong Basic and Applied Basic Research Foundation (Grant No. 2022B1515120059), the National Key R&D Program of China (Grant No. 2023YFF0725004), the Guangzhou-HKUST(GZ) Joint Funding Program (Grant No. 2023A03J0008), and the Education Bureau of Guangzhou Municipality.
@@ Appendix C Baselines
  • GNNDelete: GNNDelete freezes the model parameters and introduces a learnable deletion operator to ensure that the influence of deleted elements is removed while preserving the remaining knowledge of the model.
@@
Backbone Method PubMed CiteSeer Cora CS Physics
GraphSage Retrain 88.94±0.31 74.02±0.43 77.64±1.56 93.09±0.26 96.54±0.11
BEKM 85.86±0.22 73.63±0.62 80.70±0.87 90.49±0.28 95.34±0.11
BLPA 85.15±0.26 73.45±0.78 72.95±0.65 90.05±0.27 94.71±0.16
GIF 86.96±0.38 75.91±0.77 81.59±0.73 92.88±1.76 96.01±0.15
SR 86.58±0.28 74.04±1.50 78.41±1.12 91.31±0.44 94.88±0.14
Fast 86.70±0.32 74.97±0.78 78.60±1.11 91.18±0.52 94.93±0.07
MEGU 87.34±0.27 75.67±0.94 80.22±0.50 93.35±0.13 96.64±0.21
ETR 87.22±0.26 76.64±0.69 83.10±0.95 93.86±0.18 96.27±0.11
GIN Retrain 85.23±0.33 74.53±0.80 80.85±1.02 91.15±0.37 95.42±0.34
BEKM 83.56±0.37 68.89±3.65 78.89±2.42 87.11±0.31 93.83±0.25
BLPA 83.49±0.46 67.17±3.33 73.8±3.06 87.87±0.38 93.34±0.14
GIF 84.91±1.16 70.99±1.84 78.19±1.15 88.47±0.68 94.83±0.23
SR 84.31±0.36 68.53±2.23 79.82±1.78 89.46±0.46 92.40±0.51
Fast 84.10±0.49 67.06±2.16 80.07±1.34 89.16±0.51 93.65±0.29
MEGU 84.93±0.64 74.68±1.03 81.66±0.80 89.47±0.46 93.91±0.30
ETR 85.07±0.34 75.47±0.65 80.92±1.07 90.20±0.46 95.02±0.20
Table 8: Performance comparison in terms of model utility with GraphSage and GIN backbones.
@@
Appendix D ETR for Edge Unlearning and Feature Unlearning
$\left|D_{i}\right|\nabla_{\widehat{\omega}_{r}}L_{D_{i}})$ (36)
Appendix E Additional Experimental Results
To investigate ETR’s performance with different backbones, we conduct experiments using the GraphSage and GIN architectures, with results shown in Table 8. The results demonstrate that ETR outperforms other baselines. We also evaluate ETR’s performance on edge unlearning and feature unlearning tasks. The model’s utility, unlearning efficiency, and unlearning efficacy for edge unlearning are presented in Tables 9,11, and13, respectively. Additionally, performance on feature unlearning tasks is detailed in Tables 10,12, and14.